13. Find the numbers x and y such that the point (x, y, 1) lies on the line passing through the points (2,5,7) and (0,3,2).
Answer #1

point through two equation of line passing (m, 1Y,,ZU (1214, 2) 7-21 non y-Y 22-2 24-7 Y-YI here (, ,Y,Z) = (2,517) (22,42, Wso -2(3)+2 -12 12 쯤 S y=-2(3)+ ý = -12ts 25-12 YE 13 J point that So lies on the line passing through point (2,517) and 10, 3